3 – (x + (-4)) = 2x – 6 + 5x
Answer by checkley79(3341)

3 – (x + (-4)) = 2x – 6 + 5x
3-(x-4)=7x-6
3-x+4=7x-6
7-x-7x=-6
-8x=-6-7
-8x=-13
x=-13/-8
x=13/8 or 1 5/8 ans.